A carnival game that involves flipping a card to reveal a hidden prize costs $5.00 to play. There are two cads that reveal a $20 prize. What is the expected value for someone playing this game?
Answer by greenestamps(13208)



Technically, the question can't be answered, because you did not specify the number of cards. If the game is played with a standard deck of 52 cards, you need to say so in the statement of the problem.

I will assume the cards are a standard 52-card deck....

Use the basic definition of expected value.
           probability    value (winnings
                       minus cost of playing)
  -----------------------------------------------------------
 winning card  2/52    -5+20 = 15       (2/52)(15) =   30/52
 other card   50/52    -5              (50/52)(-5) = -250/52
                                    -------------------------
                                     expected value: -220/52

The expected value for playing the game is -220/52 dollars, or about -$4.23.